ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: 65, 42 Karoui, Abderrazek, Faculté des Sciences de Bizerte, Tunisia Wavelets: Construction and quality of approximation We construct biorthogonal wavelet filters with parametric coefficients. These filters give use to compactly supported wavelet bases of L^{2}(IR). The Sobolev regularity of the constructed wavelets is estimated by the use of a practical numerical method. Finally, we give some conditions that ensure the uniform approximation by wavelets in the Hölder spaces.